# Bayesian Shrinkage in High-Dimensional VAR Models: A Comparative Study

### v1

@@ 1 Introduction
Vector autoregressive (VAR) models remain a cornerstone of multivariate time series analysis (Sims, 1980; Chan, 2020; Koop and Korobilis, 2013). A $d$-dimensional VAR($p$) posits that an observed time series $\{\mathbf{y}_{t}\}_{t=1}^{T}$ satisfies a regression model where $\mathbf{y}_{t}$ given $p$ previous observations $\mathbf{y}_{t-p},\dots,\mathbf{y}_{t-1}$ is modeled as
$\mathbf{y}_{t}\;=\;\mathbf{A}_{1}\mathbf{y}_{t-1}\;+\;\mathbf{A}_{2}\mathbf{y}_{t-2}\;+\;\cdots\;+\;\mathbf{A}_{p}\mathbf{y}_{t-p}\;+\;\boldsymbol{\varepsilon}_{t}.$ (1)
When each $\mathbf{A}_{i}$ is a $d\times d$ coefficient matrix, and we assume that
$\boldsymbol{\varepsilon}_{t}\sim\mathcal{N}(\mathbf{0},\Sigma_{\varepsilon}),$
is normally distributed white noise, that is, $\boldsymbol{\varepsilon}_{t}$ has mean $\mathbf{0}$ and covariance $\Sigma_{\varepsilon}$ and the $\boldsymbol{\varepsilon}_{t}$’s are uncorrelated over time. VAR models see extensive use in macroeconomic data analysis, where the assumption of correlated vector time series is natural; see, for instance, Stock and Watson (2002); Crump et al. (2021); Carriero et al. (2022); Zhou and Chan (2023) for discussions on the ongoing empirical suitability and methodological refinements of VAR structures in economics.
Standard VAR theory relies on the stationarity of the time series. If the characteristic polynomial of $(I_{d}-\mathbf{A}_{1}z-\cdots-\mathbf{A}_{p}z^{p})$ does not have all its roots outside the unit circle, the process may be nonstationary or cointegrated. In such cases, one typically either differences or applies error-correction forms to restore stationarity, or else fits a specialized variant (such as a vector error-correction model).
A single VAR($p$) of dimension $d$ has $d^{2}p$ parameters, so with increasing $d$ or $p$, the parameter space quickly becomes high dimensional (Bańbura et al., 2010; Koop, 2013; Korobilis and Pettenuzzo, 2019). Consequently, regularization (often referred to as “shrinkage” in the VAR literature) is crucial not only to prevent overfitting but also to control the high variability in estimated parameters, which can undermine model stability and predictive accuracy (Stock and Watson, 2002; Huber and Feldkircher, 2017; Chan, 2021). In this context, shrinkage describes the penalization of numerous parameters toward zero or toward a simpler structure, so as to enforce parsimony. For VAR models, shrinkage is particularly needed when the number of time series variables ($d$) or the chosen lag order ($p$) is large relative to the available sample size, a common scenario in macroeconomic forecasting (Bańbura et al., 2010; Bai et al., 2022).
A variety of shrinkage techniques have been developed in both Bayesian and frequentist frameworks. Bayesian local-global priors, such as the horseshoe (Carvalho et al., 2010) and Bayesian lasso (Park and Casella, 2008), aggressively shrink small or irrelevant coefficients toward zero while allowing large signals to remain relatively unshrunk. The spike-and-slab prior (George and McCulloch, 1997) offers an alternative Bayesian framework, combining sparsity with flexibility in signal detection. Hierarchical shrinkage priors, like the horseshoe and its variants (Makalic and Schmidt, 2016; Prüser, 2021), adaptively adjust shrinkage intensity and have shown strong performance in time series applications (Prüser, 2021). Other recent Bayesian shrinkage methods in time series have been explored (Huber and Koop, 2023; Kowal et al., 2019). Frequentist approaches such as ridge regression (Doan et al., 1984) penalize large coefficients to prevent overfitting, while nonparametric shrinkage approaches can directly estimate an optimal shrinkage factor—often via James–Stein type estimators—and adjust the sample covariance-based estimates accordingly (Giannone et al., 2015; Del Negro and Giannoni, 2015).
Recent literature has refined global-local priors and explored flexible hierarchical frameworks tailored to VARs. Huber and Feldkircher (2019) propose an adaptive shrinkage scheme with a Normal-Gamma prior that learns the degree of shrinkage across coefficients, showing improved forecasts in large VAR systems. Bitto and Frühwirth-Schnatter (2019) apply a Horseshoe-type prior in time-varying parameter models, highlighting the ability of heavy-tailed priors to shrink aggressively while preserving large signals. In the realm of block-specific shrinkage, Aprigliano (2020) demonstrates that treating different coefficient groups with distinct shrinkage intensities can further boost forecast accuracy. Recent computational strategies, such as variational approximations, also facilitate faster Bayesian inference for high-dimensional VARs, as discussed in Gefang et al. (2023).
Lasso-based methods (Tibshirani, 1996) and elastic net (Zou and Hastie, 2005) also appear in the high-dimensional VAR toolkit, providing sparse solutions when many coefficients are truly zero or near zero. Recent innovations, such as dynamic shrinkage models (Griffin and Brown, 2017) and time-varying parameter (TVP) VARs (Nakajima and West, 2011), add further flexibility for capturing structural breaks or changing relationships over time.
Beyond ridge and nonparametric shrinkage, recent research has proposed more flexible or structured penalties to handle the large parameter spaces typical of VAR models. Group lasso and hierarchical lag penalties can exploit the grouped structure of lagged predictors by shrinking or eliminating entire blocks of coefficients (Nicholson et al., 2017; Basu et al., 2019; Nicholson, 2020). Non-convex penalties such as the smoothly clipped absolute deviation (SCAD) and the minimax concave penalty (MCP) balance aggressive shrinkage of small coefficients with reduced bias for large signals (Song and Bickel, 2019; Chen and Chen, 2021). Adaptive lasso variants further improve selection consistency by penalizing large coefficients less, based on preliminary estimates (Kock and Callot, 2015). Lastly, time series–specific modifications—like rolling cross-validation or penalization that accounts for serial dependence (Nicholson et al., 2017; Medeiros et al., 2021)—help tailor these methods to dynamic settings, often improving stability and forecast performance in high-dimensional VARs.
@@
In this paper, we compare five shrinkage approaches for high-dimensional VAR estimation: three Bayesian priors (horseshoe, lasso, and normal) and two frequentist estimators (ridge and nonparametric shrinkage). We evaluate each method’s handling of over-parameterization in both low- and high-dimensional settings, using root mean squared error, interval coverage, interval length, and one-step-ahead RMSE to assess parameter recovery and forecast performance.
@@
The remainder of the paper is structured as follows. Section 2 defines the VAR($p$) framework and outlines both Bayesian and frequentist shrinkage estimators. Section 3 describes our simulation designs and metrics, Section 4 presents the simulation results, and Section 5 provides the real-data application to Canadian macroeconomic variables. Finally, we offer concluding remarks.
2 Bayesian and Frequentist Approaches to VAR($p$)
@@ 2.1 Bayesian Shrinkage Priors
$\mathbf{y}_{t}\;\sim\;\mathcal{N}\!\bigl{(}\mathbf{X}_{t}\,\mathbf{B}^{\prime},\;\Sigma_{\varepsilon}\bigr{)},$
@@
where $\mathbf{L}_{\Omega}$ is the Cholesky factor of a correlation matrix with an LKJ prior (Lewandowski et al., 2009), and each component of $\sigma$ follows a half-Cauchy prior. This flexible structure permits correlation among the $d$ error components.
@@
[…]
@@
The horseshoe prior (Carvalho et al., 2010; Makalic and Schmidt, 2016) introduces more adaptive shrinkage via a local-global hierarchy. Each $\beta_{j}$ is modeled as $\beta_{j}=B_{\mathrm{raw},j}\,\lambda_{j}\,\tau$, where $B_{\mathrm{raw},j}\sim\mathcal{N}(0,1)$, $\lambda_{j}\sim\mathrm{C}^{+}(0,1)$ (local scale), and $\tau\sim\mathrm{C}^{+}(0,1)$ (global scale). Small coefficients are heavily shrunk by small local scales, while the heavy-tailed Cauchy priors allow some large signals to remain.
@@
Finally, the Bayesian lasso (Park and Casella, 2008) imposes a Laplace (double-exponential) prior,
@@ 2.2 Frequentist Methods
$\min_{\boldsymbol{\beta}\in\mathbb{R}^{d^{2}p}}\;\sum_{t=p+1}^{T}\left\lVert\mathbf{y}_{t}-\sum_{i=1}^{p}\mathbf{A}_{i}\,\mathbf{y}_{t-i}\right\rVert^{2}\;+\;\lambda\left\lVert\boldsymbol{\beta}\right\rVert_{2}^{2},$ (3)
[…]
@@
We use a James–Stein-like shrinkage approach for VAR coefficients (Giannone et al., 2015; Del Negro and Giannoni, 2015), implemented in R via VARshrink with method="ns". Instead of explicitly solving (3), the NS method estimates the necessary sample covariances of ${\mathbf{y}_{t-i},\mathbf{y}_{t}}$ and then applies a closed-form shrinkage rule to these covariance estimates, thereby producing a shrunk solution for $\bf{B}$. In this paper, we rely on the default choice for the shrinkage parameter, which VARshrink selects via a moment-based (empirical Bayes) formula akin to Stein’s unbiased risk estimate.
@@ 3.1 Data-Generating Processes
We design three VAR data scenarios to examine both low- and high-dimensional settings and the effects of overfitting the lag order. Each dataset is simulated with a 50-observation burn-in (discarded), followed by $T=200$ observations. Of these 200, the first 180 are the training set used for estimation, and the remaining 20 are held out for forecasting. Each scenario is replicated $N=50$ times with independent parameter draws of $\beta_{j}$ for each simulation.
Scenario 1 (low dimension, overfit lag).
We generate a 3-dimensional VAR(1) process here but intentionally fit a VAR(4). To form the $3\times 3$ transition matrix $A$, each of its 9 entries is set to zero with probability $0.7$ (so on average across simulations, 3 are nonzero) and the nonzeros are drawn uniformly from (0.4, 0.4)0.40.4(-0.4,\,0.4)( - 0.4 , 0.4 ). We then rescale $A$ so its largest eigenvalue in modulus is strictly below 1 by dividing by $1.1$ times the maximum absolute eigenvalue, and set the noise covariance to $\Sigma=0.05\,I_{3}$. As a result, there are $4\times 3^{2}$ parameters to estimate in B, even though only 3 of them are non-zero on average.
Scenario 2 (high dimension, correct lag).
We generate a 20-dimensional VAR(1) process and also fit a VAR(1). The $20\times 20$ matrix $A$ is constructed where each of its 400 entries is zero with probability $0.7$ (so roughly 120 are nonzero), with nonzeros drawn randomly from (0.4, 0.4)0.40.4(-0.4,\,0.4)( - 0.4 , 0.4 ). Afterward, we ensure stationarity by dividing by $1.1$ times the maximum absolute eigenvalue. We set $\Sigma=0.1\,I_{20}$.
Scenario 3 (high dimension, overfit lag).
We use the same data-generating scheme as in Scenario 2 (thus a sparse VAR(1) in 20 dimensions), but this time fit a VAR(4). As a result, there are $16004\times 20^{2}=$ 1600 parameters to estimate, even though only 120 of them are truly nonzero on average across simulations.
@@ 3.2 Estimation Methods
[…]
@@ 3.4.3 Scenario 3 (High-Dimension, Overfit Lag)
Horseshoe has the lowest parameter RMSE (0.0394), with better-than-nominal average coverage (97.5%) and intervals about 6% shorter than those of the next-best method. Lasso (0.104) and Normal (0.117) have higher RMSEs but maintain nominal coverage (95–96%) through wider intervals (0.432–0.464). Both ns and Ridge have moderate RMSEs (0.0619–0.0635) but show lower coverage (88.2% and 84.5%) and narrower intervals (0.18–0.20). Horseshoe remains the top performer in parameter RMSE for 100% of the replications.

### v2

@@ Abstract
Keywords: Vector autoregression; Bayesian shrinkage; Horseshoe prior; High-dimensional time series; Forecasting accuracy
@@ 1 Introduction
Vector autoregressive (VAR) models remain a cornerstone of multivariate time‑series analysis Sims, (1980); Chan, (2020); Koop and Korobilis, (2013). A dd‑dimensional VAR($p$) posits that an observed series $\{\mathbf{y}_{t}\}_{t=1}^{T}$ satisfies
$\mathbf{y}_{t}\;=\;\mathbf{A}_{1}\mathbf{y}_{t-1}+\mathbf{A}_{2}\mathbf{y}_{t-2}+\cdots+\mathbf{A}_{p}\mathbf{y}_{t-p}+\boldsymbol{\varepsilon}_{t},$ (1)
where $\boldsymbol{\varepsilon}_{t}\!\sim\!\mathcal{N}(\mathbf{0},\Sigma_{\varepsilon})$ is white noise. VARs are ubiquitous in macroeconomics and finance; see Stock and Watson, (2002); Crump et al., (2021); Carriero et al., (2022); Zhou and Chan, (2023) for recent surveys and methodological refinements.
A VAR becomes non‑stationary when the roots of $(I_{d}-\mathbf{A}_{1}z-\cdots-\mathbf{A}_{p}z^{p})$ lie on or inside the unit circle. Standard remedies include differencing, error‑correction representations, or Bayesian priors that down‑weight explosive parameter draws.
Because a VAR($p$) contains $d^{2}p$ coefficients, dimensionality rises quickly with either $d$ or $p$ Bańbura et al., (2010); Koop, (2013); Korobilis and Pettenuzzo, (2019). Shrinkage—penalising or down‑weighting coefficients toward simpler structures—therefore plays a critical role in stabilising estimation and forecasts Stock and Watson, (2002); Chan, (2021), especially when $d^{2}p\gg T$ Bai et al., (2022).
Bayesian shrinkage. Local–global priors such as the horseshoe Carvalho et al., (2010) or Bayesian lasso Park and Casella, (2008) shrink most coefficients strongly while allowing a few to escape toward their likelihood values. Variants include the spike‑and‑slab George and McCulloch, (1997), hierarchical horseshoes Makalic and Schmidt, (2016); Prüser, (2021), order‑invariant priors Chan et al., (2024), dynamic sparsification for hybrid TVP‑VARs Chan, (2023), and structured factor‑augmented horseshoes Zhou and Chan, (2023). Semi‑global or block‑specific shrinkage extends the idea by letting related groups share a common scale parameter Gruber and Kastner, (2025); Prüser and Blagov, (2022). Computational advances now make these priors feasible in very large systems: fast variational Bayes Bernardi et al., (2024), approximate Bayes for huge multi‑country VARs Huber et al., (2022), and conjugate subspace shrinkage Huber and Koop, (2023). Pandemic‑robust “outlier‑aware” priors improve real‑time performance during extreme events Cascaldi-Garcia, (2022). Other recent Bayesian shrinkage methods in time series have been explored Huber and Koop, (2023); Kowal et al., (2019); Katz et al., (2024, 2025).
Frequentist and machine‑learning regularisation. Ridge regression Doan et al., (1984) remains popular, while equation‑by‑equation Lasso or elastic‑net penalties Tibshirani, (1996); Zou and Hastie, (2005); Sánchez García and Cruz Rambaud, (2022) yield sparse high‑dimensional VARs with frequentist guarantees under weak dependence Masini et al., (2022). Envelope and reduced‑rank ideas provide additional dimension‑reduction tools Samadi and Herath, (2023); Cubadda and Hecq, (2022), and non‑linear factor compression has recently been explored Klieber, (2024). Structured penalties—group Lasso, hierarchical lag shrinkage, SCAD, MCP—continue to improve point forecasts and interpretability Nicholson et al., (2017); Basu et al., (2019); Nicholson, (2020); Song and Bickel, (2019); Chen and Chen, (2021). New inferential results ensure valid impulse‑response analysis even after sparse estimation Krampe et al., (2023).
Forecasting and empirical performance. Time‑varying parameter BVARs with horseshoe or normal‑gamma shrinkage adapt well to structural change Bitto and Frühwirth-Schnatter, (2019); Feldkircher et al., (2024). Factor‑adjusted network VARs (FNETS) separate common and idiosyncratic dynamics to enhance both forecasts and network interpretation Barigozzi et al., (2024). Large‑scale empirical comparisons consistently find that local–global or semi‑global priors outperform purely global shrinkage and many frequentist alternatives in density and point forecasts Huber and Feldkircher, (2019); Aprigliano, (2020); Gefang et al., (2023).
@@
In this paper, we compare five shrinkage approaches for high-dimensional VAR estimation: three Bayesian priors (horseshoe, lasso, and normal) and two frequentist estimators (ridge and nonparametric shrinkage). We evaluate each method’s handling of over-parameterization in both low- and high-dimensional settings, using root mean squared error, interval coverage, interval length, and RMSE to assess parameter recovery and forecast performance.
@@
The remainder of the paper is structured as follows. Section 2 defines the VAR($p$) framework and outlines both Bayesian and frequentist shrinkage estimators. Section 3 describes our simulation designs, metrics, and results. Section 4 provides the real-data application to Canadian macroeconomic variables. Finally, we offer concluding remarks.
[…]
@@ 2.1 Bayesian Shrinkage Priors
$\mathbf{y}_{t}\;\sim\;\mathcal{N}\!\bigl{(}\mathbf{B}\mathbf{X}_{t}^{\prime},\;\Sigma_{\varepsilon}\bigr{)},$
@@
where $\mathbf{L}_{\Omega}$ is the Cholesky factor of a correlation matrix with an LKJ prior Lewandowski et al., (2009), and each component of $\sigma$ follows a half-Cauchy prior. This flexible structure permits correlation among the $d$ error components.
@@
$\beta_{j}\;\sim\;\mathcal{N}\!\bigl{(}0,\,1\bigr{)},\quad j=1,\ldots,d^{2}p,$
@@
The horseshoe prior Carvalho et al., (2010); Makalic and Schmidt, (2016) introduces more adaptive shrinkage via a local-global hierarchy. Each $\beta_{j}$ is modeled as $\beta_{j}=B_{\mathrm{raw},j}\,\lambda_{j}\,\tau$, where $B_{\mathrm{raw},j}\sim\mathcal{N}(0,1)$, $\lambda_{j}\sim\mathrm{C}^{+}(0,1)$ (local scale), and $\tau\sim\mathrm{C}^{+}(0,1)$ (global scale). Small coefficients are heavily shrunk by small local scales, while the heavy-tailed Cauchy priors allow some large signals to remain.
@@
Finally, the Bayesian lasso Park and Casella, (2008) imposes a Laplace (double-exponential) prior,
@@
How the priors influence estimation.
The three priors differ in the shape and hierarchy of their scale parameters, and these choices translate directly into the amount and selectivity of shrinkage:
  • Normal (ridge). A single global variance forces equal shrinkage on every $\beta_{j}$, pulling all coefficients toward 0 irrespective of their signal strength. This produces stable point estimates but can underestimate posterior uncertainty when $d^{2}p\gg T$.
  • Horseshoe. Independent local scales $\lambda_{j}\!\sim\!\mathrm{C}^{+}$ give each coefficient its own amount of shrinkage, while the heavy‑tailed Cauchy hierarchy preserves large signals. The result is strong suppression of noise coefficients together with wider credible intervals for the few active ones.
  • Bayesian lasso. The exponential tails of the Laplace prior lie between the two extremes above, producing more sparsity than the Normal but heavier global shrinkage than the Horseshoe; moderate signals are therefore attenuated the most.
These qualitative differences predict the empirical patterns we later observe: ridge delivers the narrowest—but sometimes under‑covering—intervals, the horseshoe attains the best balance of low RMSE and near‑nominal coverage, and the lasso sits in between. Section 5 returns to this point in light of the simulation and data results.
@@ 2.2 Frequentist Methods
$\min_{\boldsymbol{\beta}\in\mathbb{R}^{d^{2}p}}\;\sum_{t=p+1}^{T}\left\lVert\mathbf{y}_{t}-\sum_{i=1}^{p}\mathbf{A}_{i}\,\mathbf{y}_{t-i}\right\rVert^{2}\;+\;\lambda\left\lVert\bf{B}\right\rVert_{2}^{2},$ (3)
where $\bf{B}$ is just the vectorized collection of $\{\mathbf{A}_{i}\}$. We set the regularization parameter $\lambda=.1$ in our analysis.
@@
We use a James–Stein-like shrinkage approach for VAR coefficients Giannone et al., (2015); Del Negro and Giannoni, (2015), implemented in R via VARshrink with method="ns". Instead of explicitly solving (3), the NS method estimates the necessary sample covariances of $\mathbf{y}_{t-i},\mathbf{y}_{t}$} and then applies a closed-form shrinkage rule to these covariance estimates, thereby producing a shrunk solution for $\bf{B}$. In this paper, we rely on the default choice for the shrinkage parameter, which VARshrink selects via a moment-based (empirical Bayes) formula akin to Stein’s unbiased risk estimate.
@@ 3.1 Data-Generating Processes
We examine three Monte–Carlo scenarios that differ in dimension ($d$) and in whether the fitted lag order $p$ coincides with the true lag order $p^{\star}$. Table 1 records the fixed design choices; the recipe that follows is applied independently in each of the $N_{\text{rep}}=50$ replications.
Table 1: Key parameters of each simulation scenario. “Sparsity” is the probability that an entry of $A_{1}$ is set to 0. All series use a 50-observation burn-in, followed by $T_{\mathrm{train}}=180$ training points and a $H=20$-point test set.
Scenario $d$ true $p^{\star}$ fitted $p$ sparsity $\sigma_{\varepsilon}^{2}$
1: low-$d$, over-fit 3 1 4 0.70 0.05
2: high-$d$, well-fit 20 1 1 0.70 0.10
3: high-$d$, over-fit 20 1 4 0.70 0.10
Algorithm 1 Simulation procedure for a single replication
1:Input: $(d,p^{\star},\text{sparsity},\sigma_{\varepsilon}^{2})$
2:Output: training sample $\{\mathbf{y}_{t}\}_{t=1}^{T_{\text{train}}}$ and hold-out sample $\{\mathbf{y}_{t}\}_{t=T_{\text{train}}+1}^{T_{\text{train}}+H}$
3:Draw the coefficient matrix $A_{1}$. Each entry $(A_{1})_{ij}$ is independently
$(A_{1})_{ij}\;=\;\begin{cases}0,&\text{with probability }\text{sparsity},\\[4.0pt] U(-0.4,0.4),&\text{otherwise.}\end{cases}$
All higher-order true lags are set to zero: $A_{2}=A_{3}=\dots=A_{p^{\star}}=\mathbf{0}$.
4:Stationarity margin. Compute the spectral radius $\rho_{\max}$ of $A_{1}$ and rescale $A_{1}\leftarrow A_{1}/(1.1\,\rho_{\max})$ so that all roots of $I_{d}-A_{1}z$ are strictly inside the unit circle.
5:Simulate the innovations. Draw $\boldsymbol{\varepsilon}_{t}\sim\mathcal{N}(\mathbf{0},\sigma_{\varepsilon}^{2}I_{d})$ independently for $t=-49,\dots,T_{\text{train}}+H$.
6:Generate the series. Starting from $\mathbf{y}_{-50}=\mathbf{0}$, evolve $\mathbf{y}_{t}=A_{1}\mathbf{y}_{t-1}+\boldsymbol{\varepsilon}_{t}$.
7:Burn-in. Discard the first 50 observations ($t=-49,\dots,0$); the retained series therefore begins at $t=1$.
8:Split the data. Keep $t=1,\dots,T_{\text{train}}$ for estimation and $t=T_{\text{train}}+1,\dots,T_{\text{train}}+H$ for forecast evaluation.
Stationarity margin.
Following the textbook discussion in Lütkepohl, (2005, Sec. 2.1), we scale the simulated companion matrix so that its spectral radius is comfortably inside the unit circle. Specifically, we divide the initial draw $A_{1}$ by $1.1\,\rho_{\max}$, where $\rho_{\max}$ is the largest eigenvalue (in modulus). This simple rescaling keeps all roots of the characteristic polynomial strictly below one—thereby ensuring covariance–stationarity—while preserving empirically plausible coefficient magnitudes for macro‑VAR applications.
@@ 3.2 Estimation Methods
We fit the three Bayesian models by calling Stan with 4 parallel Markov chains, each run for 2000 total iterations (the first 500 of which are warm-up). We fix seed=123 for reproducibility and use $\{\texttt{adapt\_delta}=0.9,\;\texttt{max\_treedepth}=12\}.$ In Stan’s Hamiltonian Monte Carlo (HMC) framework, adapt_delta is the target acceptance probability, and increasing it to 0.9 aims for smaller step sizes and more conservative sampling. The max_treedepth parameter caps the depth of the binary tree in each iteration’s leapfrog integrator, preventing extremely long trajectories.
@@ 3.4.3 Scenario 3 (High-Dimension, Overfit Lag)
Horseshoe has the lowest parameter RMSE (0.0394), with better-than-nominal average coverage (97.5%) and intervals about 8% shorter than those of the next-best method. Lasso (0.104) and Normal (0.117) have higher RMSEs but maintain nominal coverage (95–96%) through wider intervals (0.432–0.464). Both ns and Ridge have moderate RMSEs (0.0619–0.0635) but show lower coverage (88.2% and 84.5%) and narrower intervals (0.182–0.197). Horseshoe remains the top performer in parameter RMSE for 100% of the replications.